1.In a factory, 7 workers get a monthly salary of rs. 1400 each and other 8 workers get a monthly salary of rs. 1600 each. Find the avg salary of a worker.
2. The avg score of 10 players is 42. If each player makes 12 runs extra in a match ,what will be their new avg.?

Answer by Cromlix(4381)

7 * 1400 = 9800
8 * 1600 = 12800
Total = 22600
For average, divide by total number of workers (15)
rs 22600/15 = rs 1506.67.
..................
If avg score of 10 players = 42
then total = 10 * 42 = 420
If each player makes 12 extra runs = 120
420 + 120 = 540
540/10 = 54 (The new average)
Hope this helps.
:-)